# Weize Chen1  , Jiarui Yuan1∗, Chen Qian1Cheng Yang2 Zhiyuan Liu1, Maosong Sun1 1

### v1

@@ 3.1 Benchmark Results
Table 1 showcases Optima’s performance across a diverse set of tasks, revealing consistent improvements over baseline methods in both effectiveness and efficiency. In IE tasks, Optima variants demonstrate substantial gains, particularly in multi-hop reasoning scenarios like HotpotQA and 2WMHQA. Here, iSFT-DPO achieves peak performance while significantly reducing token usage compared to the strongest baseline SC. Notably, on 2WMHQA, iSFT-DPO improves F1 score by 38.3% (2.8x improvement) while using only 10% of the tokens required by MAD. This trend extends to other information exchange tasks, where Optima variants maintain high performance with drastically lower token counts. The debate tasks present a more nuanced picture, yet Optima’s benefits remain evident. Better task performance and token efficiency are still observed in ARC-C and MMLU, but for the MATH and GSM8k tasks, Optima variants show comparable or slightly lower performance than SC, but still with much higher token efficiency. We conjecture this is due to the task’s difficulty and the small size of their training set. However, as we will demonstrate in Section 3.2, Optima models trained on MATH transfer effectively to GSM8k, achieving performance nearly equivalent to models trained directly on GSM8k, with high token efficiency. More interestingly, Section 3.3 will show that applying SC to Optima variants trained on MATH or GSM8k leads to better inference scaling laws on GSM8k compared to CoT SC.
A closer look at Optima variants reveals interesting trade-offs. Optima-iSFT often prioritizes performance at the expense of token efficiency, demonstrating the poorest efficiency in 5 of 8 tasks. In contrast, Optima-iDPO often achieves remarkable reductions in token usage, occasionally with performance trade-offs. Optima-iSFT-DPO emerges as a robust compromise, frequently delivering top-tier performance with satisfying token efficiency.
3.2 How Well Does Optima Generalize to OOD Tasks?
@@
These results underscore Optima’s potential for developing adaptable MAS, demonstrating that Optima-trained models learn transferable skills for efficient information exchange and collaborative reasoning. However, transferring to more distant domains remains challenging, e.g., we find it hard to transfer from HotpotQA to CBT, or from MATH to ARC-C. We believe it is a promising area for future research to explore if scaling Optima to more generalized multi-task training could enhance the generalization of communication strategies in LLMs.
3.3 Can Optima lead to Better Inference Scaling Law?
(a) Inference scaling on debate tasks
(b) Performance vs. token usage on GSM8k
@@ 3.3 Can Optima lead to Better Inference Scaling Law?
Figure 3: Optima’s impact on inference scaling laws. (a) Relationship between Optima variants’ self-consistency steps and performance on debate tasks. Solid lines represent majority voting accuracy, while dashed lines show coverage. (b) Performance of various models on GSM8k as a function of token usage, demonstrating Optima’s efficiency gains.
@@
Recent research has highlighted the importance of inference scaling laws, which describe how model performance improves with increased compute during inference, typically by generating multiple samples per problem (Brown et al., 2024; Wu et al., 2024). While training scaling laws focus on the relationship between model size, dataset size, and performance, inference scaling laws explore the trade-off between inference compute budget and task accuracy. This paradigm offers a promising avenue for enhancing model capabilities without the need for further training models.
@@
Fig. 3 illustrates Optima’s impact on inference scaling laws. The left panel shows the relationship between the number of SC steps and performance on multi-agent debate tasks. We observe that while majority voting accuracy tends to plateau after a certain number of steps, the coverage, defined as the percentage of problems answered correctly at least once, continues to improve logarithmically with increased sampling. This trend aligns with findings in recent inference scaling law studies (Wu et al., 2024; Chen et al., 2024a) and suggests that more sophisticated answer selection techniques could further boost Optima’s performance. We provide additional scaling law figures for all Optima variants and on both IE and debate tasks in Appendix A, where similar trends can be observed.
@@
The right panel of Fig. 3 demonstrates Optima’s efficiency in improving inference scaling laws on the GSM8k task. Optima variants, both those trained directly on GSM8k and those transferred from MATH, consistently outperform the CoT SC baseline except the iSFT variant transferred from MATH. Notably, iDPO trained on GSM8k achieves the performance of CoT-SC at around 10,000 tokens with 88.5% fewer tokens, effectively “shifting the curve left”. This significant reduction in token usage translates to substantial computational savings without sacrificing accuracy. Moreover, the MATH-trained Optima variants, except iSFT, also deliver better inference scaling laws on GSM8k compared with CoT SC, underscoring the framework’s ability to generalize effectively across related tasks.
@@
These results highlight Optima’s potential to reshape inference scaling laws for LLM-based MAS and even general LLM systems. By enabling more efficient use of the inference compute budget, Optima allows for better performance at lower computational costs or higher performance at the same cost. This efficiency gain opens new possibilities for leveraging advanced inference techniques like weighted voting or best-of-N selection (Wu et al., 2024), potentially leading to even greater performance improvements.
@@
3.4 How Does Optima Evolve Agent Communication and Performance?
To understand the impact of different components in our reward function, we conducted an ablation study on two representative tasks: 2WMHQA for IE and ARC-C for debate. We examined the performance of Optima variants by removing either the token count regularization (#Tokens) or the LM loss (Loss) from the reward function. The results aim to answer two key questions: (1) How does token count regularization affect the efficiency-performance trade-off? (2) What is the role of language modeling loss in maintaining communication quality? Our findings consistently demonstrate the crucial role of each reward component in balancing task performance, communication efficiency, and language quality.
Table 3 presents the results of our ablation study. Removing the token count led to a substantial increase in the number of generated tokens across settings, with a particularly pronounced effect in the debate task. While this increased verbosity occasionally resulted in marginal performance improvements, it came at a significant computational cost. Conversely, eliminating the LM loss resulted in a decrease in token usage, often producing the most concise outputs among all variants. Examples comparing communication with and without LM loss can be found in Appendix C. Without LM loss, the model often generated overly concise messages containing insufficient information and was prone to hallucination, potentially explaining the inferior performance under this condition. These results underscore that effective LLM-based MAS should optimize not only for task performance but also for the efficiency and quality of inter-agent dialogue. The design of Optima’s reward function enables this holistic optimization, leading to more effective and efficient multi-agent collaboration while highlighting the delicate balance required in optimizing such systems.
@@ 3.5 How Agent Communication Evolves over Optimization Iterations?
Fig. 1 illustrates the performance gains and token efficiency of Optima variants across the optimization iterations, revealing a distinctive two-phase optimization pattern. In the initial phase (iterations 0-1), we observe a substantial improvement in task performance for all Optima variants, accompanied by a clear increase in token usage. This suggests that Optima initially prioritizes effectiveness, allowing agents to develop sophisticated problem-solving strategies through expanded communication. The subsequent iterations demonstrate Optima’s ability to refine these strategies for efficiency without compromising performance. We observe a gradual but consistent decrease in token usage across all variants, coupled with continued performance improvements.
To provide concrete examples of how Optima shapes agent communication, we present a case from iSFT on an information exchange task in Fig. 4. The base model exhibits unfocused and repetitive exchanges, failing to efficiently address the task at hand. At iteration 0, while more structured, the exchange is verbose and includes unnecessary metadata. By iteration 2, we observe a marked shift towards concise, task-oriented communication, with agents adopting a streamlined format that efficiently conveys key information. The final iteration demonstrates further refinement, with agents maintaining the efficient structure while eliminating any residual verbosity. This progression aligns with our quantitative findings, showcasing Optima’s ability to form communication patterns that are both highly effective and remarkably efficient.
@@ 4 Related Work
LLM-Based MAS. LLM-based MAS have emerged as a powerful paradigm for addressing complex tasks across various domains. Seminal works by Liang et al. (2023) and Du et al. (2024) demonstrated the potential of LLM-powered agents in collaborative problem-solving through multi-agent debate. This foundation has sparked diverse research directions, including role-playing for complex reasoning (Wang et al., 2024b; Chen et al., 2024b), collaborative software development (Qian et al., 2024c; Hong et al., 2024; Ishibashi & Nishimura, 2024), and embodied agent interactions (Zhang et al., 2024; Mandi et al., 2024; Guo et al., 2024). Recent studies have shown that increasing the number and diversity of agents can lead to performance gains in MAS (Wang et al., 2024a; Li et al., 2024a; Chen et al., 2024c). However, as LLM-based MAS grow in scale and complexity, challenges related to computational costs and communication efficiency become more pronounced (Chen et al., 2024d; Li et al., 2024b). Notably, there is a lack of systematic training algorithms specifically designed to optimize both the effectiveness and efficiency of LLM-based multi-agent systems, with most existing approaches relying on updating agent memory (Qian et al., 2024a; Gao et al., 2024). Our work addresses this gap by introducing a training framework that simultaneously enhances communication efficiency and task effectiveness in LLM-based MAS.
@@ 5 Conclusion
We present Optima, a novel framework for training LLM-based MAS that significantly improves communication efficiency and task performance. Extensive experiments across a range of tasks demonstrate Optima’s consistent superiority over both single-agent and multi-agent baselines. The framework introduces key innovations such as iterative training techniques, a balanced reward function, and an MCTS-inspired approach for data generation. Optima also shows promise in enhancing inference scaling laws and transferring knowledge to OOD tasks. These findings highlight the critical role of efficient communication in MAS and LLM systems. While Optima marks a major step forward in multi-agent LLM training, further exploration into its scalability to larger models and more complex scenarios is a promising direction for future research.
@@ Appendix B Additional Pseudo-Codes for Optima Variants
To elucidate the implementation of various Optima variants, we present algorithmic representations of several critical processes intrinsic to these variants. Specifically, we delineate the pseudo-code for (1) the initialization dataset collection process, as elucidated in Section 2.2 and illustrated in Algorithm 3; (2) the Monte Carlo Tree Search-based data generation process employed in iDPO (Section 2.4), as depicted in Algorithm 5; and (3) the procedure for node selection during the expansion phase of MCTS, as outlined in Algorithm 4. These algorithmic representations serve to provide a comprehensive and rigorous exposition of the methodological framework underlying the Optima variants.
@@ Appendix C Case Study on Reward Components Ablation
[”Film”, ”Last Of The Long-Haired Boys”],
[”Release”, ”1968”]
@@
[”Film”, ”Age Of The Dragons”],
[”Release”, ”2011”]
@@
Appendix D Case Study on Debate Task
@@
In Section 3.5, we presented an example from 2WMH QA, illustrating Optima’s impact on an information exchange task. Here, we provide a complementary case study from a debate task to demonstrate Optima’s effectiveness across different multi-agent settings. Fig. 6 showcases the evolution of agent communication in a debate task across iterations 0, 2, and 4 of Optima training. The task involves discussing the environmental impact of fertilizer runoff on ocean bays.
@@
Appendix E Experiment Details
E.1 Data Generation

### v2

@@ 3.1 Benchmark Results
Table 1 showcases Optima’s performance across diverse tasks, revealing consistent improvements in effectiveness and efficiency. For IE tasks, Optima variants excel, particularly in multi-hop reasoning like HotpotQA and 2WMHQA. iSFT-DPO achieves the best performance while significantly reducing token usage compared to SC. Notably, on 2WMHQA, iSFT-DPO improves F1 by 38.3% (2.8x) while using only 10% of MAD’s tokens. This efficiency extends to other IE tasks, where Optima variants maintain high performance with drastically lower token counts.
In debate tasks, Optima’s benefits are nuanced but evident. It achieves better performance and efficiency on ARC-C and MMLU, while on MATH and GSM8k, Optima variants show comparable or slightly lower performance than SC, but with much higher token efficiency. We attribute this to task difficulty and limited training data. Nevertheless, Section 3.2 will show Optima models trained on MATH transfer effectively to GSM8k, achieving near-equivalent performance with high efficiency. Additionally, Section 3.3 will demonstrate that applying SC to Optima variants trained on MATH or GSM8k greatly improves inference scaling laws on GSM8k compared to CoT SC.
Among Optima variants, iSFT often prioritizes performance at the cost of efficiency, while iDPO achieves remarkable token reductions, sometimes with performance trade-offs. iSFT-DPO strikes a robust balance, frequently delivering top-tier performance with satisfying efficiency. Results on Llama 3.2 3B in Appendix F further validate Optima’s robustness.
3.2 How Well Does Optima Generalize?
@@
These results underscore Optima’s potential for developing adaptable MAS, demonstrating that Optima-trained models learn transferable skills for efficient information exchange and collaborative reasoning. However, transferring to more distant domains remains challenging, e.g., we find it hard to transfer from from MATH to ARC-C. We believe it is a promising area for future research to explore if scaling Optima to more generalized multi-task training could enhance the generalization.
3.3 Can Optima Improve Inference Scaling?
@@
Figure 3: Optima’s impact on inference scaling laws. Left Relationship between Optima variants’ self-consistency steps and performance on debate tasks. Solid lines represent majority voting accuracy, while dashed lines show coverage. Right Performance of various models on GSM8k as a function of token usage, demonstrating Optima’s efficiency gains.
@@
Recent research emphasizes inference-time scaling, which describes how model performance improves with increased compute during inference, typically by generating multiple samples per problem (Brown et al., 2024; Wu et al., 2024). Unlike training scaling laws, which focus on model size, dataset size, and performance, inference-time scaling explore the trade-off between compute budget and task accuracy, offering a promising way to enhance model capabilities without further training.
@@
Fig. 3 illustrates Optima’s impact on inference-time scaling. The left panel shows the relationship between SC steps and performance on multi-agent debate tasks. While majority voting accuracy plateaus after a certain number of steps, coverage (the percentage of problems answered correctly at least once) improves logarithmically with increased sampling. This aligns with recent studies (Wu et al., 2024; Chen et al., 2024a), suggesting advanced answer selection techniques could further boost Optima’s performance. Additional scaling law figures for all Optima variants and tasks are in Appendix A, where similar trends are observed.
@@
The right panel demonstrates Optima’s efficiency in improving inference scaling laws on GSM8k. Optima variants, including those transferred from MATH, consistently outperform CoT SC, except for MATH-trained iSFT. Notably, GSM8k-trained iDPO matches CoT-SC performance with 88.5% fewer tokens, effectively “shifting the curve left". This reduction in token usage translates to significant computational savings without sacrificing accuracy. MATH-trained Optima variants, except iSFT, also deliver better scaling laws on GSM8k than CoT SC, highlighting Optima’s cross-task generalization.
@@
These results underscore Optima’s potential to reshape inference-time scaling for MAS and general LLM systems. By enabling more efficient use of compute budgets, Optima achieves better performance at lower costs or higher performance at the same cost. This efficiency opens possibilities for integrating advanced inference techniques like weighted voting or tree-search (Wu et al., 2024), potentially leading to further performance gains.
3.4 How Does Optima Evolve Performance?
To understand the impact of reward function components in our reward function, we conducted an ablation study on 2WMHQA (IE) and ARC-C (debate). We removed either token count regularization (#Tokens) or LM loss (Loss) to address: (1) How does token count regularization affect efficiency-performance trade-offs? (2) What role does LM loss play in maintaining communication quality? Our findings highlight the importance of each component in balancing performance, efficiency, and language quality.
Table 3 presents the results of our ablation study. Removing the token count leads to a substantial increase in the number of generated tokens across settings, with a particularly pronounced effect in the debate task. While this increased verbosity occasionally results in marginal performance improvements, it comes at a significant computational cost. Conversely, eliminating the LM loss results in a decrease in token usage, often producing the most concise outputs among all variants. Examples comparing communication with and without LM loss can be found in Appendix C. Without LM loss, the model often generates overly concise messages containing insufficient information and is prone to hallucination, potentially explaining the inferior performance. Overall, Optima’s reward function achieves the balance among task effectiveness, token efficiency and dialogue quality, enabling effective and efficient multi-agent collaboration.
@@ 3.5 How Agent Communication Evolves over Optimization Iterations?
Fig. 1 shows the performance gains and token efficiency of Optima variants across optimization iterations, revealing a two-phase pattern. In the initial phase (iterations 0-1), all variants show significant performance improvements alongside increased token usage, indicating Optima prioritizes effectiveness by enabling agents to develop sophisticated strategies through expanded communication. In later iterations, Optima refines these strategies for efficiency without sacrificing performance, with token usage decreasing gradually while performance continues to improve.
Concrete examples of Optima’s impact on agent communication are provided in Appendix D (iSFT on an information exchange task) and Appendix E (debate task). The base model tends to produce verbose, repetitive exchanges, while Optima-trained models exhibit more concise and task-oriented communication.
3.6 Can Optima Scale with More Agents?
While the previous experiments highlight Optima’s effectiveness in two-agent scenarios, which is a controlled setting that circumvents issues such as communication order and effectively validates the framework, we also evaluate its scalability in three-agent settings for IE and debate tasks. The results, detailed in Appendix G, demonstrate that Optima continues to enhance both effectiveness and efficiency.
@@ 4 Related Work
LLM-Based MAS. LLM-powered multi-agent systems have demonstrated success in collaborative problem-solving through approaches like multi-agent debate (Liang et al., 2023; Du et al., 2024). Subsequent work explores role-playing for reasoning (Wang et al., 2024b; Chen et al., 2024b), software development (Qian et al., 2024c; Hong et al., 2024), and embodied interactions (Zhang et al., 2024; Mandi et al., 2024), with scale and diversity improving performance (Wang et al., 2024a; Li et al., 2024a). However, efficiency challenges emerge as systems grow (Chen et al., 2024c; Qian et al., 2024d), with existing methods focusing on memory updates rather than comprehensive training (Qian et al., 2024a). Our framework addresses this gap through joint optimization of communication efficiency and task effectiveness.
@@
Inference-Time Scaling and Token Efficiency. Compute scaling has enhanced LLM capabilities, with approaches like majority voting and reward-guided tree search improving performance on reasoning tasks (Chen et al., 2024a; Wu et al., 2024; Brown et al., 2024; Saad-Falcon et al., 2024). However, these methods increase computational demands, highlighting the need for token efficiency. Recent work achieves efficiency through latent space reasoning via step distillation (Deng et al., 2023, 2024; Hao et al., 2024; Cheng and Durme, 2024), but at the cost of comprehensibility. Our framework advances this by (1) demonstrating iterative training framework that improves both token efficiency and task effectiveness in MAS context, and (2) showing that enhanced efficiency enables more sampling within fixed compute budgets, leading to better inference-time scaling.
@@ 5 Conclusion
We introduce Optima, a novel framework for training LLM-based MAS that significantly enhances communication efficiency and task performance. Experiments show Optima’s consistent superiority over single-agent and multi-agent baselines. We introduce key innovations such as iterative training, a balanced reward function, and MCTS-inspired data generation. Crucially, Optima effectively improves inference-time scaling and transfers effectively to OOD tasks, underscoring the importance of efficient communication in MAS and LLM systems. While Optima marks a major step forward in multi-agent LLM training, further exploration into its scalability to larger models and more complex scenarios is a promising direction for future research.
Limitations
While Optima demonstrates significant improvements in communication efficiency and task effectiveness for LLM-based multi-agent systems, our study has several limitations. First, our experiments primarily focus on two-agent scenarios with a shared model architecture, leaving open questions about scaling to larger teams (e.g., 5-10 agents) and heterogeneous agent configurations. Although preliminary results with three agents show promising trends (Section 3.6), the dynamics of larger groups may introduce new challenges in coordination efficiency that require further investigation. Second, while we demonstrate cross-task generalization within similar domains (e.g., MATH to GSM8k), transferring Optima-trained models to substantially different application areas (e.g., from QA to math or coding) remains unexplored. Finally, while we evaluate on standard benchmarks, real-world deployment scenarios may involve additional constraints that our framework does not explicitly address. These limitations highlight valuable directions for future research rather than fundamental flaws, as Optima’s core contributions, iterative optimization with efficiency-aware rewards and MCTS-inspired data generation, provide a flexible foundation adaptable to these extensions.
@@ Appendix A Inference Scaling Laws on Information Exchange Tasks
Algorithm 2 Iterative Supervised Fine-Tuning
2:Optimized model $\mathcal{M}_{T}$
4:for $t=0$ to $T-1$ do
5:     $\mathcal{D}_{t}^{*}\leftarrow\emptyset$
6:     for each $d_{i}\in\mathcal{D}$ do
9:         if $R(\tau_{i}^{*})>\theta_{\text{sft}}$ then
10:              $\mathcal{D}_{t}^{*}\leftarrow\mathcal{D}_{t}^{*}\cup\{(d_{i},\tau_{i}^{*})\}$
11:         end if
12:     end for
15:end for
16:return $\mathcal{M}_{T}$
Algorithm 3 Iterative Direct Preference Optimization
2:Optimized model $\mathcal{M}_{T}$
4:for $t=0$ to $T-1$ do
5:     $\mathcal{D}_{t}^{\text{DPO}}\leftarrow\emptyset$
6:     for each $d_{i}\in\mathcal{D}$ do
9:     end for
11:end for
12:return $\mathcal{M}_{T}$
@@ Appendix B Additional Pseudo-Codes for Optima Variants
To elucidate the implementation of various Optima variants, we present algorithmic representations of several critical processes intrinsic to these variants. Specifically, we delineate the pseudo-code for (1) the initialization dataset collection process, as elucidated in Section 2.2 and illustrated in Algorithm 1; (2) the iterative supervised fine-tuning process introduced in Section 2.3 and shown in Algorithm 2; (3) the iteratiove DPO process as detailed in Section 2.4 and illustrated in Algorithm 3; (4) the Monte Carlo Tree Search-based data generation process employed in iDPO (Section 2.4), as depicted in Algorithm 5; and (5) the procedure for node selection during the expansion phase of MCTS, as outlined in Algorithm 4. These algorithmic representations serve to provide a comprehensive and rigorous exposition of the methodological framework underlying the Optima variants.
@@ Appendix C Case Study on Reward Components Ablation
["Film", "Last Of The Long-Haired Boys"],
["Release", "1968"]
@@
["Film", "Age Of The Dragons"],
["Release", "2011"]
@@
Appendix D Case Study on Information Exchange Task
In this section, we present a case study from iSFT on an information exchange task, with the evolution of agent communication detailed in Fig. 5.
The base model exhibits unfocused and repetitive exchanges, failing to efficiently address the task at hand. At iteration 0, while more structured, the exchange is verbose and includes unnecessary metadata. By iteration 2, we observe a marked shift towards concise, task-oriented communication, with agents adopting a streamlined format that efficiently conveys key information. The final iteration demonstrates further refinement, with agents maintaining the efficient structure while eliminating any residual verbosity. This progression aligns with our quantitative findings, showcasing Optima’s ability to form communication patterns that are both highly effective and remarkably efficient.
Appendix E Case Study on Debate Task
@@
In Appendix D, we presented an example from 2WMH QA, illustrating Optima’s impact on an information exchange task. Here, we provide a complementary case study from a debate task to demonstrate Optima’s effectiveness across different multi-agent settings. Fig. 6 showcases the evolution of agent communication in a debate task across iterations 0, 2, and 4 of Optima training. The task involves discussing the environmental impact of fertilizer runoff on ocean bays.
@@
Appendix F Results on Llama 3.2 3B
Information Exchange Debate
HotpotQA 2WMH QA MATH GSM8k ARC-C
Method F1 #Tok F1 #Tok Acc #Tok Acc #Tok Acc #Tok
CoT 22.7 355.8 16.5 235.0 46.3 556.7 78.7 288.9 51.5 256.1
SC ($n=8$) 28.0 2804.6 24.2 467.7 56.8 4436.0 88.6 2300.4 57.6 2068.6
MAD 31.8 1677.9 27.6 2152.8 46.3 2509.2 81.2 763.8 37.4 872.4
AutoForm 22.8 87.6 19.9 106.5 42.7 629.2 77.6 443.9 22.9 265.9
Optima-iSFT 53.2 54 65.2 47.7 46.1 585.4 81.8 313.9 62.7 156.2
Optima-iDPO 49.4 59.9 57.0 65.4 47.4 575.7 81.4 290.8 63.1 132.7
Optima-iSFT-DPO 52.5 48.7 66.8 51.4 46.8 548.4 80.8 270.1 61.6 141.4
Table 6: the results with the base model being Llama 3.2 3B
As illustrated in Section 3.1, to verify Optima’s ability of generalizing to different base models, we conduct experiment based on Llama 3.2 3B. The results are presented in Table 6. From the results, we can see that Optima is still able to significantly improve both efficiency and performance for the model with smaller parameter sizes.
Appendix G Results on Scenarios with More Agents
2WMH QA ARC-C
Setting F1 #Tok Acc #Tok
CoT 20.5 139.8 65.2 138.9
SC(n=8) 28.7 1052.8 75.6 1116.7
MAD(2-agent) 25.9 543.7 71.4 478.0
AutoForm 22.6 147.8 59.1 128.2
iSFT 62.0 62.8 72.6 123
iDPO 56.3 55.8 75.6 76.2
iSFT-DPO 60.7 53.7 75.4 72.7
Table 7: the results on three-agent scenarios
Table 7 presents the results on three-agent scenarios. We select one task from both the IE task and the debate task for experimentation. It is important to note that in the debate task, we no longer designate a specific agent as the solver and another as the critic, which differs from the two-agent scenarios.
In the IE task, the 3-agent setting generally performs worse than the 2-agent setting due to the more distributed nature of the information, but Optima still offers performance gain against baselines. In the debate task, Optima also continues to provide a performance boost while significantly reducing token usage.
Appendix H Experiment Details
H.1 Data Generation